OpenStudy (anonymous):

Solve the equation 25 = radical 5^(2x + 6). Using complete sentences, explain the strategy used in solving this equation.

OpenStudy (phi):

I assume this is \[25=\sqrt[5]{2x+6}\] ? if so, you raise the radical to the 5th power, to "undo it" of course, you have to do the same thing to both sides of the equation so \[25^5=\left(\sqrt[5]{2x+6}\right)^5 \\ \]

OpenStudy (raden):

is your equation like this ?